A flute has a length of 58.0cm. If the speed of sound in air is 343m/s, what is the fundamental frequency of the flute, assuming it is a tube closed at one end and open at the other?


1
Expert's answer
2021-02-17T12:08:13-0500

Answer

For fundamental frequency

"\lambda=4L\\=4(0.58) =2.32m"



And fundamental frequency will be

"f=\frac{v}{\lambda}=\frac{343}{2.32}\\f=147.8HZ"